11 8 15 12 11 12 11 14 15 5 6 11

16 14 13 14 13 9 11 19 12 10 13 5

13 7 3 14 13 16 20 15 5 15 15 5

15 7 9 11 17 18 10 14 9 7 4 15

13 12 16 8 12 12


(a)  Find:

i.              Mean

ii.             Median

iii.            Mode

iv.            Interquartile range

v.             Range

vi.            Variance

vii. Standard deviation


1
Expert's answer
2021-07-22T12:11:26-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 X & f \\ \hline
 3 & 1 \\
 \hdashline
 4 & 1\\
 \hdashline
 5 & 4\\
 \hdashline
 6 & 1\\
 \hdashline
 7 & 3\\
 \hdashline
 8 & 2\\
 \hdashline
 9 & 3\\
 \hdashline
 10 & 2\\
 \hdashline
 11 & 6 \\
 \hdashline
 12 & 6\\
 \hdashline
 13 & 6\\
 \hdashline
 14 & 5\\
 \hdashline
 15 & 7\\
 \hdashline
 16 & 3\\
 \hdashline
 17 & 1\\
 \hdashline
 18 & 1\\
 \hdashline
 19 & 1\\
 \hdashline
 20 & 1\\
 \hdashline
 & \sum f=54\\
 \hdashline
\end{array}"

i.


"mean=\bar{X}=\dfrac{1}{n}\displaystyle\sum_{i=1}^nX_i=\dfrac{630}{54}=\dfrac{35}{3}\approx11.666667"


ii.


"X_{27}=12=X_{28}"

"median=12"

iii.

"mode=15,"

appeared 7 times.


iv.

Lower quartile "(X_L):9"


Upper quartile "(X_U):15"


Interquartile range

"IQR=X_U-X_L=15-9=6"



v.


"Range=20-3=17"

vi.


"Var(X)=\sigma^2=\dfrac{1}{n}\displaystyle\sum_{i=1}^n(X_i-\bar{X})^2=\dfrac{830}{54}=\dfrac{415}{27}"

"\approx15.370370"

vii.


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{415}{27}}\approx3.920506"